OpenStudy (anonymous):

Find the equation of a line in slope-intercept form, using the point slope form then graph 2x-3/2y=-3

OpenStudy (anonymous):

@zepdrix @DangerousJesse

OpenStudy (xapproachesinfinity):

miss worded question! do you need that equation of a line in slope intercept form and then graph it you question make no sense, read it again to see what im talking about

OpenStudy (jdoe0001):

hmmm.... so.you're supposed to make it a point-slope form firstly? and then a slope-intercept one?

OpenStudy (jdoe0001):

you could also just take a quick screenshot of the material and post it here

OpenStudy (anonymous):

this is what my homework says exactly :/

OpenStudy (xapproachesinfinity):

i don't thin so? do you have it in textbook

OpenStudy (anonymous):

im going to assume i need to put it into y=mx+b then graph it?

OpenStudy (xapproachesinfinity):

well that is what i think your question is about! but check pls what it is exactly

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

we haven't been assigned textbooks yet, this is just a work sheet it the very top it says Find the equation of a line in slope-intercept form, using the point slope form. then, this individual question says graph 2x-3/2y=-3 but its all under the same set of instructions

OpenStudy (jdoe0001):

you could also just take a quick screenshot of the material and post it here :)

OpenStudy (xapproachesinfinity):

yeah screenshot the whole thing and post it^_^

OpenStudy (anonymous):

ok, i have to scan it first since im on my computer, one sec

OpenStudy (jdoe0001):

hmmm is on paper?

OpenStudy (jdoe0001):

you can also just snapshot it with a webcam.... surely you have one or a cellphone camera

OpenStudy (jdoe0001):

just make sure is clear and legible =)

OpenStudy (xapproachesinfinity):

if you have in your computer just click the prt sc button it should be copied then paste it into word document and post then

OpenStudy (xapproachesinfinity):

if you have it on computer*

OpenStudy (anonymous):

well this is physical paper, its not online

OpenStudy (xapproachesinfinity):

Okay, then do as he said. use phone camera

OpenStudy (anonymous):

its number 15

OpenStudy (jdoe0001):

hmmm 15 is just graphing only

OpenStudy (xapproachesinfinity):

that one is just about graphing the line 2x-3/2y=-3

OpenStudy (anonymous):

but i only know how to graph it when its in y=mx+b

OpenStudy (jdoe0001):

yeap is only \(\bf graph\qquad 2x-\cfrac{3}{2}y=-3\)

OpenStudy (xapproachesinfinity):

choose two points to graph it the easier choise is x=0 y=? y=0 x=?

OpenStudy (anonymous):

im going to delete that just cause it has my whole name so write it down if you haven't!

OpenStudy (anonymous):

y=0?

OpenStudy (xapproachesinfinity):

yeah it's okay you can delete

OpenStudy (xapproachesinfinity):

iif y=0 what is x=? if x=0 what is y=?

OpenStudy (xapproachesinfinity):

plug x=0 into the equation

OpenStudy (jdoe0001):

at least you don't have your email so we can spam you with bargains on steak knives with a free garden hose =)

OpenStudy (anonymous):

XD

OpenStudy (xapproachesinfinity):

\(\Large \rm\color{midnightblue}{2(0)-\frac{3}{2}y=-3}\) solve for y

OpenStudy (xapproachesinfinity):

lol

OpenStudy (anonymous):

i got -3/-3/2=y :[

OpenStudy (anonymous):

that doesnt sound right

OpenStudy (xapproachesinfinity):

no that's fine you just need to simplify that fraction

OpenStudy (xapproachesinfinity):